OpenStudy (anonymous):

Solve for v = (17 + x)v = 25g

OpenStudy (anonymous):

can you rewrite this question plz

OpenStudy (anonymous):

simple- divide both sides by \((17+x)\):\[(17 + x)v = 25g\]divide:\[{{(17 + x)v}\over(17+x)} = {{25g}\over(17+x)}\]\[{{\cancel{(17 + x)}v}\over\cancel{(17+x)}} = {{25g}\over(17+x)}\]\[v={{25g}\over(17+x)}\]

OpenStudy (anonymous):

I thought so. Thanks

OpenStudy (anonymous):

the objective for solving for such variables is literally just getting them alone on one side...so just see what you can do to get it alone....YW :)